A ball thrown vertically upward returns to its starting point at 4s. Find its initial speed.


1
Expert's answer
2022-05-20T04:53:41-0400

The total time of motion

"t=2\frac{v_0}{g}"

Hence

"v_0=gt/2\\
=9.8*4/2=19.6\:\rm m/s"